a particle moves along a path y=2cos3x in such a way that the component of its velocity in the x-direction is 4 units. find its acceleration?
1
Expert's answer
2019-12-03T10:43:08-0500

The condition of the problem regarding the x-direction means "x^{'}_t=4". We can immediately conclude "x^{''}_{tt}=0" and "x=x_0+4\cdot{t}", where "x_o=x(t=0)". This is uniform movement on x. Substitute this expression in y we find "y=2cos(3x_0+12t)".

Thus velocity in the y-direction is

"y^{'}_t=-2\cdot{12}\cdot{sin(3x_0+12t)}=-24sin(3x_0+12t)"

and aceleration

"y^{''}_{tt}=-24\cdot{12}\cdot{cos(3x_0+12t)}=-288cos(3x_0+12t)"

Answer: Acceleration along the axis x is 0. The acceleration a particle in y-direction "-288cos(3x_0+12t)" where "x_0=x(t=0)" particle position at the initial moment of time.